The upper and lower specifications for a service are 11.5 min. and 9.0 min., respectively. The process average is 9.50 min. If 6-sigma process capability is desired:

(a) what should the process mean be? (Provide your answer rounded to 2 decimal places.)

(b) what should the process standard deviation be? (Provide your answer rounded to 4 decimal places.)
Answer #1

Here, Given Data,

Upper Specification Limit or USL = 11.5 Minutes

Lower Specification Limit or LSL = 9.0 Minutes

Current Process Mean = 9.50 Minutes,

As 6 Sigma Process Capability is Desired, Thus,

Process Capability Cp has to be 2,

Cp = 2 = (USL-LSL)/(6* Standard Deviation),

Or, (11.5-9.0)/( 6* Standard Deviation) = 2

Or, Standard Deviation = 2.5/12 = 0.2083 Minutes

Lower Bound for the Process Capability Index or Cpl has to be 2

Cpl = 2 = (Process Mean – LSL)/3*Standard Deviation,

Or, (Process Mean – 9.0)/3*0.2083=2,

Or, Process Mean = 9.0 + 1.25 = 10.25 Minutes

Upper Bound for the Process Capability Index or Cpu has to be 2 or

(USL – Process Mean)/3*Standard Deviation= 2,

(11.5 – Process Mean)/(3*0.2083) = 2,

Or, Process Mean = 11.5 – 1.25 = 10.25 Minutes,

Answer for a)

Hence, Process Mean Should be 10.25 Minutes,

Answer for b)

Hence, Process standard Deviation should be 0.2083 Minutes